OpenStudy (anonymous):

Let A_0=1 and a_1=1 use generating functions to solve the recurrence equation a_(n+2)= a_(n+1)+ 2a_n for n>=0

OpenStudy (anonymous):

so i tried to work this out and i started with my G(x)=a_o+a_1x+a_2X^2+... -1X(G(x) -2X^2 G(x) and got (1-x-2X^2) G(x) = 1-2x^2 G(x)= 1-2x^2 / (1-x-x^2) but i am not sure if it right or where to go from there... and help would be great thanks :)

OpenStudy (amistre64):

i was considering a homogenous and aprticular solution similar to diffy Qs

OpenStudy (amistre64):

a_(n+2)= a_(n+1)+ 2a_n a_(n+2) - a_(n+1) - 2a_n = 0 for some reason, the subs can go exponential r^(n+2) - r^(n+1) - 2r^n = 0 r^n (r^2 - r - 2) = 0 when r = whatever the qf gives us i spose

OpenStudy (amistre64):

r = -1, 2 .... but then it gets fuzzy again

OpenStudy (zarkon):

\[a_n=C2^n+D(-1)^n\] use \(a_0=1,a_1=1\) to find \(C,D\)

myininaya (myininaya):

I think this is a nice recipe http://www.cs.columbia.edu/~cs4205/files/CM2.pdf

myininaya (myininaya):

You have to do some manipulation so you can use \[A(x)=\sum_{n=0}^{\infty}a_nx^n \] Like at one point I was using: \[\sum_{n=0}^{\infty}a_{n+2}x^{n}=x^{-2}\sum_{n=0}^{\infty}a_{n+2}x^{n+2}=x^{-2}(A(x)-a_2x^2-a_1x^1-a_0)\]

myininaya (myininaya):

Or You can call big A big G. whatever

OpenStudy (amistre64):

a 67 page recipe eh ... not quite rachel ray approved

OpenStudy (amistre64):

if my head was not about to explode id love to read thru it better :)

myininaya (myininaya):

page 10-11

myininaya (myininaya):

I made a error

myininaya (myininaya):

\[\sum_{n=0}^{\infty}a_{n+2}x^{n}=x^{-2}\sum_{n=0}^{\infty}a_{n+2}x^{n+2}=x^{-2}(A(x)-a_1x^1-a_0) \]